xcoledd1:

How do you find the slant height for a pyramid if your is height is 9.1 cm and all 4 sides is 12 cm.

xcoledd1:

@acehuert

xcoledd1:

i just need the slant height

acehuert:

15.06

xcoledd1:

0_0

xcoledd1:

thx

acehuert:

np

xcoledd1:

closing the post

wio:

 This is just the Pythagorean theorem. If you take a look at the cross section, one of the legs is the height \(9.1cm\), and the other leg is half the side, or \(12cm/2\). The slant height is just the hypotenuse:\[ x^2 = (9.1cm)^2+(12cm/2)^2\implies x = \sqrt{9.1^2+6^2}cm =\sqrt{118.81}cm =10.9cm \]